OpenStudy (anonymous):

A fireman has to reach a burning building. Determine the length of the shortest ladder that will reach over a 2 metre high fence to the burning building which is 1 metre behind the fence. Need the 2 Equations to sub so I can then take derivative and set to 0! someone please find my 2 equations I have the length of ladder one but need the limiting equation!!!

OpenStudy (anonymous):

z² = x² + y²

OpenStudy (anonymous):

yes, and now the limiting equation? that I will use to substitute into c^2 = a^2 + b^2 to obtain a single variable on the right side and then do the derivative

OpenStudy (anonymous):

the length of ladder is limited!

OpenStudy (anonymous):

=> xdx/dt + ydy/dt = 0

OpenStudy (anonymous):

do I need to carry the exponent on Z over before deriving?

OpenStudy (anonymous):

It's constant, so "No"

OpenStudy (anonymous):

ok so I have \[L^2 = (y+1)^2 + (x+2)^2\]

OpenStudy (anonymous):

Do you post the whole problem at all?

OpenStudy (anonymous):

yea thats the whole thing

OpenStudy (anonymous):

see what I mean I need a second equation to limit that one to one variable before I can derive

OpenStudy (anonymous):

Is it Related Rate problem? because without the data of dx/dt and dy/dt, it doesn't look right!

OpenStudy (anonymous):

no its a optimization problem from grade 12 calculus... part of a take home assignment

OpenStudy (anonymous):

Let me check, but keep Z to take derivative first!

OpenStudy (anonymous):

Do you how to apply similar triangles?

OpenStudy (anonymous):

xy = 2 => y = 2/x That's limitation!

OpenStudy (anonymous):

OK i had that before, so I am subbing that into the original hypotenuse equation the deriving and setting to zero?

OpenStudy (anonymous):

then*

OpenStudy (anonymous):

After taking derivative, subs y = 2/x, solve for x, y, z

OpenStudy (anonymous):

so take the derivative of L2=(y+1)2+(x+2)2 then sub?

OpenStudy (anonymous):

sorry the 2s are exponents

OpenStudy (anonymous):

?

OpenStudy (bahrom7893):

wait I'm not following, I gotta start from the beginning